# High-dimensional Autoregressive Modeling for Time Series with Hierarchical Structures

### v1

@@ 4.2 Hyperparameter selection
Some hyperparameters involved in Algorithm 1 are typically unknown in most real-world applications, and their selection not only impacts the model’s performance, but also reveals the underlying structural information of the data. This subsection introduces a boosting-based procedure for determining the number of action orders $K_{y}$ and $K_{x}$, the corresponding action sets $\mathcal{P}_{y}(N,K_{y})$ and $\mathcal{P}_{x}(M,K_{x})$, and the associated rank sequences $\{\boldsymbol{s}_{k^{\prime}}\}_{k^{\prime}=1}^{K_{y}}$ and $\{\boldsymbol{r}_{k}\}_{k=1}^{K_{x}}$.
To guide the selection of these hyperparameters, we adopt an information criterion. Let $\mbox{$\mathscr{\widehat{A}}$}(R)$ denote the estimator from (3.1), where the configuration $R$ collects all hyperparameters: $R=(K_{y},K_{x},\mathcal{P}_{y}(N,K_{y}),\mathcal{P}_{x}(M,K_{x}),\{\boldsymbol{s}_{k^{\prime}}\}_{k^{\prime}=1}^{K_{y}},\{\boldsymbol{r}_{k}\}_{k=1}^{K_{x}})$. The Bayesian information criterion (BIC) is then constructed below,
$\mathrm{BIC}(R)=T\log\left\{\mathcal{L}_{T}(\mbox{$\mathscr{\widehat{A}}$}(R))\right\}+\phi d(R)\log(T),$ (4.3)
where $\phi$ is a tuning parameter, and $d(R)=\sum_{k^{\prime}=1}^{K_{y}}\sum_{n=1}^{N}q_{n}s_{n-1,k^{\prime}}s_{n,k^{\prime}}+\sum_{k=1}^{K_{x}}\sum_{n=1}^{N}q_{n}r_{n-1,k}r_{n,k}+Ls_{\mathrm{tot}}r_{\mathrm{tot}}$ is the total number of parameters.
The boosting-based hyperparameter selection procedure consists of two main stages: action order selection and rank reduction, which are detailed below. The overall procedure is summarized in Algorithm 2. Finally, the lag length $L$ is determined via a simple grid search to minimize the BIC over a predetermined candidate set $\{1,2,\ldots,L_{\max}\}$.
4.2.1 Action order selection
To construct the boosting-based algorithm, we first introduce the key notations and setup. Define $\widetilde{\boldsymbol{\Lambda}}_{x}(\alpha_{k},\boldsymbol{r}_{k})=\boldsymbol{\Lambda}_{x}(\{\alpha_{k}\})$, where the operator $\boldsymbol{\Lambda}_{x}(\cdot)$ is defined in (2.10) and $\boldsymbol{r}_{k}$ is the associated rank. It can be verified that for any two positive integer vectors $\boldsymbol{r}_{k,1}$ and $\boldsymbol{r}_{k,2}$, there exists a matrix $\widetilde{\boldsymbol{\Lambda}}_{x}(\alpha_{k},\boldsymbol{r}_{k,3})$ with $\boldsymbol{r}_{k,3}\leq\boldsymbol{r}_{k,1}+\boldsymbol{r}_{k,2}$ by elementwise comparison, satisfying
$\widetilde{\boldsymbol{\Lambda}}_{x}(\alpha_{k},\boldsymbol{r}_{k,3})=\boldsymbol{O}\left[\widetilde{\boldsymbol{\Lambda}}_{x}(\alpha_{k},\boldsymbol{r}_{k,1}),\widetilde{\boldsymbol{\Lambda}}_{x}(\alpha_{k},\boldsymbol{r}_{k,2})\right],$ (4.4)
where $\boldsymbol{O}$ is an invertible matrix that can be absorbed into $\boldsymbol{\Theta}$. The same applies to $\widetilde{\boldsymbol{\Lambda}}_{y}(\cdot)$ by defining $\widetilde{\boldsymbol{\Lambda}}_{y}(\beta_{k^{\prime}},\boldsymbol{s}_{k^{\prime}})=\boldsymbol{\Lambda}_{y}(\{\beta_{k^{\prime}}\})$, where $\boldsymbol{s}_{k^{\prime}}$ is the associated rank and $\boldsymbol{\Lambda}_{y}(\cdot)$ is defined similarly to (2.10). The proof is provided in Supplementary Material.
We refer to any model defined by (4.1) with $\boldsymbol{\Lambda}_{y}=\widetilde{\boldsymbol{\Lambda}}_{y}(\beta_{k^{\prime}},\boldsymbol{1}_{N})$ and $\boldsymbol{\Lambda}_{x}=\widetilde{\boldsymbol{\Lambda}}_{x}(\alpha_{k},\boldsymbol{1}_{N})$ as a weak learner, denoted by $WL(\beta_{k^{\prime}},\alpha_{k})$, where $\boldsymbol{1}_{N}\in\mathbb{R}^{N}$ is a vector of ones. This weak learner serves as the base component in our boosting framework. The additive property introduced in (4.4) implies that the proposed model is additive with respect to rank increments: increasing the rank of a selected action order is equivalent to concatenating additional weak learners that share the same action order.
To reduce computational complexity, we predefine candidate sets of action orders, denoted by $\mathcal{P}_{y}^{(0)}$ and $\mathcal{P}_{x}^{(0)}$, based on prior knowledge of the data. These sets contain all feasible action orders to be considered for responses and predictors, respectively. For notational convenience, we abuse $K_{y}$ and $K_{x}$ to denote the number of action orders in $\mathcal{P}_{y}^{(0)}$ and $\mathcal{P}_{x}^{(0)}$ in this section.
To track how many times each action order is selected during the boosting process, we define the evolving action sets at iteration $v$ as $\mathcal{P}_{y}^{(v)}=\{\beta_{1}(s_{1}^{(v)}),\ldots,\beta_{K_{y}}(s_{K_{y}}^{(v)})\}$ and $\mathcal{P}_{x}^{(v)}=\{\alpha_{1}(r_{1}^{(v)}),\ldots,\alpha_{K_{x}}(r_{K_{x}}^{(v)})\}$, where $s_{k^{\prime}}^{(v)}$ and $r_{k}^{(v)}$ indicate the number of times action orders $\beta_{k^{\prime}}$ and $\alpha_{k}$ have been selected up to iteration $v$. We set $\mathcal{P}_{y}^{(0)}=\{\beta_{1}(0),\ldots,\beta_{K_{y}}(0)\}$ and $\mathcal{P}_{x}^{(0)}=\{\alpha_{1}(0),\ldots,\alpha_{K_{x}}(0)\}$ for initialization. For each selected action order $\beta_{k^{\prime}}(s_{k^{\prime}}^{(v)})\in\mathcal{P}_{y}^{(v)}$ and $\alpha_{k}(r_{k}^{(v)})\in\mathcal{P}_{x}^{(v)}$, we let $\boldsymbol{s}_{k^{\prime}}^{(v)}=s_{k^{\prime}}^{(v)}\boldsymbol{1}_{N}$ and $\boldsymbol{r}_{k}^{(v)}=r_{k}^{(v)}\boldsymbol{1}_{N}$. These sets and ranks are then used in Algorithm 1 to estimate $\boldsymbol{\Lambda}_{y}^{(v)}$, $\boldsymbol{\Lambda}_{x}^{(v)}$ and $\boldsymbol{\Theta}^{(v)}$, yielding the prediction $\widehat{\boldsymbol{y}}_{t}^{(v)}$.
@@ 4.2.1 Action order selection
$s_{\tilde{k}^{\prime}}^{(v+1)}=s_{\tilde{k}^{\prime}}^{(v)}+1,\quad r_{\tilde{k}}^{(v+1)}=r_{\tilde{k}}^{(v)}+1$
while keeping all other counts unchanged. By iteratively adding weak learners structured according to selected action orders, the model progressively refines its representational capacity in a data-adaptive and interpretable way.
4.2.2 Rank reduction
Algorithm 2 Boosting-Based Algorithm for Hyperparameter Selection
Input: Data $\{(\boldsymbol{y}_{t},\boldsymbol{x}_{t})\}_{t=1}^{T}$, initial sets $\mathcal{P}_{y}^{(0)}$ and $\mathcal{P}_{x}^{(0)}$, and maximum iterations $V_{\max}$.
Initialize: Set $\widehat{\boldsymbol{y}}_{t}^{(0)}=\boldsymbol{0}$ for all $t$ and $v=0$.
repeat
  Evaluate all candidate action order pairs $(\beta_{k^{\prime}},\alpha_{k})\in\{\beta_{1},\ldots,\beta_{K_{y}^{0}}\}\times\{\alpha_{1},\ldots,\alpha_{K_{x}^{0}}\}$:
  for all $(\beta_{k^{\prime}},\alpha_{k})$ do
   Fit weak learner $WL(\beta_{k^{\prime}},\alpha_{k})$ to regress $\boldsymbol{y}_{t}-\widehat{\boldsymbol{y}}_{t}^{(v)}$ on $\boldsymbol{x}_{t}$.
  end for

### v2

@@ 4.2 Hyperparameter selection
We first consider the case that the sets of action orders $\mathcal{P}_{x}(M,K_{x})$ and $\mathcal{P}_{y}(N,K_{y})$ are known, and their cardinalities, $K_{x}$ and $K_{y}$, are small; see, for example, the number of all possible action orders is not too large. A validation method can then be used to select the ranks $\{r_{m,k},1\leq m\leq M,1\leq k\leq K_{x}\}$ and $\{s_{n,k^{\prime}},1\leq n\leq N,1\leq k^{\prime}\leq K_{y}\}$.
Specifically, for an observed time series $\{\mbox{$\mathscr{Y}$}_{t},1-L\leq t\leq T\}$, we set the training and validation sets to be the first $T_{0}+L$ and last $T-T_{0}$ observations, respectively. A rolling forecasting method with the fixed starting point is then employed, and we will choose the ranks with minimum validation errors. Note that, for an action order from $\mathcal{P}_{x}(M,K_{x})$ or $\mathcal{P}_{y}(N,K_{y})$, its associated ranks will be set to zero if it is not involved in extracting features. Moreover, it is impossible to search among all possible ranks since there are many interim ranks. We suggest selecting these ranks sequentially, i.e., we choose a rank among $1\leq r_{m,k}\leq R_{\max}$ or $1\leq s_{n,k^{\prime}}\leq S_{\max}$, while all the other ranks are kept unchanged. This will depend on the order of ranks, and the selected ranks are also not optimal globally. Practically, we may iterate the procedure several times to alleviate this effect to some extent.
We next introduce a boosting method to select the sets of action orders, $\mathcal{P}_{x}(M,K_{x})$ and $\mathcal{P}_{y}(N,K_{y})$, as well as the maximum possible ranks, $R_{\max}$ and $S_{\max}$, in a data-driven manner. To this end, the Bayesian information criterion (BIC) is given below,
$\mathrm{BIC}=T\log\left\{\mathcal{L}_{T}(\mbox{$\mathscr{\widehat{A}}$})\right\}+d_{\mathrm{AR}}\log(T),$
where $d_{\mathrm{AR}}=\sum_{k=1}^{K_{x}}d_{x}(\alpha_{k})+\sum_{k=1}^{K_{y}}d_{y}(\beta_{k})+Lsr$ is the model complexity. Moreover, for an action order $\alpha=(\alpha_{(1)},\alpha_{(2)},\ldots,\alpha_{(M)})\in\mathcal{P}_{x}(M)$ with ranks $\boldsymbol{r}=(r_{1},\ldots,r_{M})$, we denote $\boldsymbol{\Lambda}_{x,\alpha}(\boldsymbol{r})=\boldsymbol{T}_{x}^{\top}(\alpha)(\boldsymbol{I}_{p_{\alpha_{(2)}}\cdots p_{\alpha_{(M)}}}\otimes\boldsymbol{G}_{1})\cdots(\boldsymbol{I}_{p_{\alpha_{(M)}}}\otimes\boldsymbol{G}_{M-1})\boldsymbol{G}_{M}\in\mathbb{R}^{P\times r_{M}}$; see (2.10) for more details. Consider the action order $\alpha$ with two groups of ranks, $\boldsymbol{r}_{1}$ and $\boldsymbol{r}_{2}$, and it is equivalent to the case with ranks $\boldsymbol{r}_{3}=\boldsymbol{r}_{1}+\boldsymbol{r}_{2}$, up to a linear transformation; see the proposition below.
Proposition 2.
Consider two vectors of ranks $\boldsymbol{r}_{1}=(r_{1,1},\ldots,r_{M,1})$ and $\boldsymbol{r}_{2}=(r_{1,2},\ldots,r_{M,2})$, and there exist a matrix $\boldsymbol{O}\in\mathbb{R}^{r_{M,3}\times(r_{M,1}+r_{M,2})}$ and the corresponding component matrices such that ${\boldsymbol{\Lambda}}_{x,\alpha}(\boldsymbol{r}_{3})=\boldsymbol{O}({\boldsymbol{\Lambda}}_{x,\alpha}(\boldsymbol{r}_{1}),{\boldsymbol{\Lambda}}_{x,\alpha}(\boldsymbol{r}_{2}))$, where $\boldsymbol{r}_{3}=\boldsymbol{r}_{1}+\boldsymbol{r}_{2}=(r_{1,3},\ldots,r_{M,3})$.
The results are similar to $\boldsymbol{\Lambda}_{y}$. Let $\mathcal{P}_{x}(M,K_{x})=\{\alpha_{1}(i_{1}),\ldots,\alpha_{K_{x}}(i_{K_{x}})\}$ and $\mathcal{P}_{y}(N,K_{y})=\{\beta_{1}(j_{1}),\ldots,\beta_{K_{y}}(j_{K_{y}})\}$ be the running sets of action orders, where $\alpha_{k}(i_{k})$ refers to the action order $\alpha_{k}$ with ranks $(r_{1,k},\ldots,r_{M,k})=(i_{k},\ldots,i_{k})$, and $\beta_{k}(j_{k})$ to the action order $\beta_{k}$ with ranks $(s_{1,k},\ldots,s_{N,k})=(j_{k},\ldots,j_{k})$. We next introduce a procedure to update $\mathcal{P}_{x}(M,K_{x})$ and $\mathcal{P}_{y}(N,K_{y})$ by considering action orders with all ranks being one:
  • (i.)
    Fit a tensor autoregressive model at (2.12) and (2.13) with the running sets of $\mathcal{P}_{x}(M,K_{x})$ and $\mathcal{P}_{y}(N,K_{y})$ to the observed time series, and denote the fitted values by $\{\widehat{\mbox{$\mathscr{Y}$}}_{t}\}$.
  • (ii.)
    Consider a regression model with the response and predictors being $\mbox{$\mathscr{Y}$}_{t}-\widehat{\mbox{$\mathscr{Y}$}}_{t}$ and $\mbox{$\mathscr{Y}$}_{t-1},\cdots,\mbox{$\mathscr{Y}$}_{t-L}$, respectively, and we try each possible action order $\alpha\in\mathcal{P}_{x}(M)$ with ranks $(r_{1},\ldots,r_{M})=(1,\ldots,1)$ and $\beta\in\mathcal{P}_{y}(N)$ with ranks $(s_{1},\ldots,s_{N})=(1,\ldots,1)$. Denote by $\alpha^{*}$ and $\beta^{*}$ the action orders with minimum mean square errors.
  • (iii.)
    Update the sets of $\mathcal{P}_{x}(M,K_{x})$ and $\mathcal{P}_{y}(N,K_{y})$. Specifically, we add $\alpha^{*}(1)$ to $\mathcal{P}_{x}(M,K_{x})$ and replace $K_{x}$ by $K_{x}+1$ if $\alpha^{*}$ is a new action order. Otherwise, we simply replace $\alpha_{k}(i_{k})$ by $\alpha_{k}(i_{k}+1)$ if $\alpha^{*}=\alpha_{k}$. $\mathcal{P}_{y}(N,K_{y})$ can be updated in the same way.
The above procedure stops when the BIC no longer decreases, and it ends up with the sets of action orders, as well as the associated maximum ranks.
5 Simulation studies
This section conducts two simulation experiments to evaluate action order misspecification, and to check finite-sample performance of the proposed methodology, respectively.
The first experiment is to evaluate the impact of action order misspecification. Consider a true action order $\alpha^{*}$ with loading matrices $\{\boldsymbol{G}_{m}^{*}\}_{m=1}^{M}$. For another action order $\alpha_{k}$, by Proposition 1, there exists a set of loading matrices $\{\boldsymbol{G}_{m,k}\}_{m=1}^{M}$ such that the feature extraction mappings under both action orders are equivalent, i.e., $\mathcal{M}(\mathcal{T}(\mbox{$\mathscr{X}$}_{t};\alpha^{*});\{\boldsymbol{G}_{m}^{*}\}_{m=1}^{M})=\mathcal{M}(\mathcal{T}(\mbox{$\mathscr{X}$}_{t};\alpha_{k});\{\boldsymbol{G}_{m,k}\}_{m=1}^{M})$, but it may lead to larger ranks and hence large model complexity when $\alpha_{k}\neq\alpha^{*}$. To numerically verify this point, we fix $M=3$ and specify the tensor dimensions as $(p_{1},p_{2},p_{3})=(8,9,10)$. Note that there are six possible action orders in total, and the true action order is given by $\alpha^{*}=(1,2,3)$, with true ranks $r_{1}^{*}=r_{2}^{*}=r_{3}^{*}=2$. The data are generated by $\boldsymbol{f}_{t}=\mathcal{M}(\mathcal{T}(\mbox{$\mathscr{X}$}_{t};\alpha^{*});\{\boldsymbol{G}_{m}^{*}\}_{m=1}^{3})\in\mathbb{R}^{r_{3}^{*}}$, where entries of $\mbox{$\mathscr{X}$}_{t}\in\mathbb{R}^{p_{1}\times p_{2}\times p_{3}}$ are $i.i.d.$ standard normal, and $\boldsymbol{G}_{m}^{*}$’s come from QR decomposition on random matrices with $i.i.d$ standard normal entries. We then directly regress $\boldsymbol{f}_{t}$ on $\mbox{$\mathscr{X}$}_{t}$, leading to a model without noise and $\boldsymbol{\Lambda}_{y}$ and $\boldsymbol{\Theta}$ being fixed as identity matrices. The sample size is set to $T=1000$, and we compare the accuracy of feature extraction with each action order $\alpha_{k}$, where $\alpha_{1}=\alpha^{*}$. When fitting the models with different action orders, we fix $r_{1,k}=r_{2,k}=r_{k}$ and $r_{3,k}=2$ for all $k$ for simplicity. Algorithm 1 is applied to estimate $\boldsymbol{G}_{m,k}$’s, and the rank $r_{k}$ varies from 2 to 8. Figure 2(a) gives mean squared errors (MSE), averaged over 50 replications.
@@ 5 Simulation studies
It can be seen that, for the true action order $\alpha_{1}$, the MSE is zero when $r_{k}$ is as small as two, and this is as expected since the model is correctly specified without noise. Moreover, for the misspecified action orders, the MSEs are substantially higher when $r_{k}$ is small, but decrease gradually to zero as $r_{k}$ increases. This is consistent with theoretical results at Proposition 1. On one hand, it confirms that loading matrices with different action orders can represent the same feature extraction mapping, given sufficiently large ranks. On the other hand, since the model complexity $d=\sum_{i=1}^{3}p_{i}r_{i}^{2}$ grows quadratically with respect to $r_{k}$, the misspecification may lead to larger model complexity and computational burden.